Perpendicular to the graph of 5x-3y=9 and passes through the orgin
Answer by mananth(16946)

5x-3y=9
3y=5x-9
/3
y=(5/3)x-3
slope = (5/3)
passing through (0,0) and perpendicular to the above line
so slope will be negative reciprocal = (-3/5)
The equation is
(y-0)=(-3/5)(x-0)
y= -3/5 x